# Meta results on data-driven control of nonlinear systems

### v1

@@ Abstract
This note aims to provide a systematic understanding of direct data-driven control, enriching the existing literature not by adding another isolated result, but rather by offering a comprehensive, versatile, and unifying framework that sets the stage for future explorations and applications in this domain. To this end, we formulate the nonlinear design problem from a high-level perspective as a set of desired controlled systems and propose systematic procedures to synthesize data-driven control algorithms that meet the design requirements specified in the desired set. Various examples are presented to demonstrate the comprehensiveness and adaptability of the proposed approach.
@@ I Introduction
The expanding interest in data-driven methodologies is fuelled by their potential to circumvent the limitations of model-based techniques, particularly in the realm of nonlinear systems where a priori given mathematical model may not be available. Nonlinear systems, characterized by their rich dynamics that are not amenable to linear analysis and necessitate a more involved approach to control design. As such, the field has seen diverse data-guided strategies being proposed and tested, ranging from virtual reference control [1], kernel-based design [2], intelligent PID [3] and sampled-data model-free control [4, 5] to data-enabled model predictive control [6], dynamic mode decomposition [8], Koopman design [7], and data-based semidefinite programs [9, Subsection V.B], [10].
@@
The current work belongs to the category of direct data-driven control design aiming to reduce the task of controller design to programs that are stated in terms of data collected from the system. Due to its inherent complexity, the landscape of direct data-driven control for nonlinear systems is fragmented with solutions that often target specific problem settings or system types. Examples include stabilization of bilinear systems [11, 12], polynomial systems [13, 14], rational systems [15], and flat systems [16]. Such specialization is valuable as it provides concrete solutions to the control problems under investigation. Nevertheless, there is a compelling need to develop a coherent framework that can address a broader spectrum of control challenges. While the focus on particular instances or system models has driven substantial contributions, a unified approach could illuminate the underlying principles, fostering even greater advancements in the field.
@@
Recognizing the need and the benefits for such a unified approach, this note introduces a meta-framework for a direct data-driven design of nonlinear control systems. We, first, embed the design specifications in a set of desired closed-loop systems. Subsequently, we establish systematic procedures to obtain the control gain from the system data such that the design specifications are enforced. The controller parametrization in [9, 17], serves as a pivotal element of the meta procedures laid out in this manuscript. The proposed approach distinguishes itself by its comprehensiveness and unifying aspects, aiming not only to incorporate several existing results into a cohesive framework but also to extend their applicability to obtain new results in the domain of nonlinear direct data-driven control. Furthermore, the proposed results provide a systematic treatment of deriving control algorithms from data. This systematic procedure ensures that diverse design objectives, such as stabilization, performance, and input-output properties can be seamlessly adapted and integrated into the design.
@@
The structure of the manuscript is as follows. The problem of interest is formulated in Section II. The solution to this problem is provided in Subsection III-A. Necessity of a key assumption is discussed in Subsection III-B. Extensions of the results to open systems are provided in Subsection III-C. Finally, the manuscript closes with conclusions in Section IV.
@@ II Problem setup
$x^{+}=f(x)+Bu$
where $x\in\mathbb{R}^{n}$, $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is locally Lipschitz, and $B\in\mathbb{R}^{n\times m}$. The notation $x^{+}$ denotes either the state variables shifted in time or the time derivative of the states depending on the setting (i.e. continuous or discrete time). While the map $f$ and the input matrix $B$ are assumed to be unknown, we assume that there exists a known library of nonlinear functions, denoted by $Z:\mathbb{R}^{n}\rightarrow\mathbb{R}^{s}$, $s\geq n$, such that $f$ can be written as $AZ(x)$ for some matrix $A\in\mathbb{R}^{n\times s}$. Hence, the nonlinear system is represented by
@@
Under a nonlinear state feedback protocol $u=KZ(x)$, $K\in\mathbb{R}^{m\times s}$, the closed-loop system admits the form
$x^{+}=F_{K}Z(x)$
with $F_{K}:=A+BK$, $F_{K}\in\mathbb{R}^{n\times s}$. The aim of the nonlinear controller is to impose some desired stability/performance properties on the closed-loop system. Rather than fixing a particular choice of a closed-loop, we consider a set of desired closed-loop dynamics. A set of desired closed-loop systems can be parametrized as
@@
where $\mathcal{V}$ denotes the set of Lyapunov functions satisfying
$\alpha\left\|x\right\|^{2}\leq V(x)\leq\beta\left\|x\right\|^{2}$
for some $\alpha,\beta>0$. Restricting to quadratic Lyapunov functions, i.e. $V(x)=x^{T}Px$, the latter can be written more explicitly as
@@
Clearly, in the above examples, any $F^{\star}\in\mathcal{F}_{\rm des}$ is such that the closed-loop system (2) meets the design criterion. $\square$
The central problem of this note is formulated below.
@@
Clearly, not any desired closed-loop system is attainable. Noting (1), a closed-loop system $x^{+}=FZ(x)$ is attainable if and only if there exists $K$ such that $A+BK=F$. Hence, the set of attainable, not necessarily desired, closed-loop system can be parametrized by the following set:
Notice that Problem 1 is solvable if and only if $\mathcal{F}_{\rm des}\cap\mathcal{F}_{\rm att}^{\rm model\,}\neq\emptyset$.
@@ III-A Solutions to Problem 1
$X_{1}=AZ_{0}+BU_{0}.$ (4)
The equality (4) is consistent with (1); specifically, $U_{0}$ corresponds to the input samples, $Z_{0}$ corresponds to the library $Z(\cdot)$ evaluated at the state samples, and $X_{1}$ stores the shifted in time (or time-derivative) state samples. Note that working with either the model or the data does not affect the choice or formulation of $\mathcal{F}_{\rm des}$. On the other hand, since we aim to use the collected data as a proxy for the model, we require a counterpart of (3) in terms of data. To this end, we define
@@
Proof. Item (i): If $\mathcal{F}_{\rm att}^{\rm data\,}$ is nonempty, then $\operatorname{im}I_{s}\subseteq\operatorname{im}Z_{0}$ and thus $Z_{0}$ has full row rank. Conversely, if $Z_{0}$ has full row rank, then $X_{1}Z_{0}^{+}\in\mathcal{F}_{\rm att}^{\rm data\,}$ for any right inverse $Z_{0}^{+}$ of $Z_{0}$. Item (ii): Observe that the set $\mathcal{F}_{\rm att}^{\rm model\,}$ in (3) can be equivalently written as
@@
where we used (4) to write the first equality. Item (iii): Bearing in mind that $\left[\begin{smallmatrix}A&B\\ I_{s}&0\end{smallmatrix}\right]$ has full column rank, the inclusion in (7) can be replaced by equality if and only if the matrix $\left[\begin{smallmatrix}Z_{0}\\ U_{0}\end{smallmatrix}\right]$ has full row rank. $\blacksquare$
Recall that Problem 1 is solvable if and only if $\mathcal{F}_{\rm att}^{\rm model\,}\cap\mathcal{F}_{\rm des}\neq\emptyset$; namely, the desired set $\mathcal{F}_{\rm des}$ and the attainable set $\mathcal{F}_{\rm att}^{\rm model\,}$ should at least share a common element. Motivated by the fact that we work directly with the data rather than the model, we replace $\mathcal{F}_{\rm att}^{\rm model\,}$ by its data-based subspace $\mathcal{F}_{\rm att}^{\rm data\,}$ in the latter solvability condition, which results in the following assumption:
@@
Proof. The fact that the set $\mathcal{K}$ is nonemempty follows from Assumption 1 and (5). Now, suppose $K\in\mathcal{K}$. Then, we find that
@@
where the second and last equality follow from (9), and the third one from (4). As $F^{\star}\in\mathcal{F}_{\rm des}$, we conclude that $(A+BK)\in\mathcal{F}_{\rm des}$ and thus the controller $KZ(x)$ solves Problem 1. $\blacksquare$
Theorem 3 provides a meta-procedure for solving Problem 1 using data:
@@
Next, we apply the above procedure to a couple of examples.
@@
As the first example, we consider the control objective of rendering the equilibrium asymptotically stable by stabilizing the linearized dynamics. It is easy to see that the set of desired closed-systems takes the form:
@@
where $Z^{\prime}$ is shorthand notation for the Jacobian of $Z$. Following the second step of the procedure, we look for $F^{\star}\in\mathcal{F}_{\rm des}\cap\mathcal{F}_{\rm att}^{\rm data\,}$. This gives rise to the following set of constraints:
$FZ^{\prime}(0)P$ $(FZ^{\prime}(0))^{T}-P<0,\quad P>0,$
$F$ $=X_{1}G,$
$I_{s}$ $=Z_{0}G,$
The aforementioned constraints can be transformed into a LMI through standard change of variables, specifically by defining $Y:=GZ^{\prime}(0)P$ and employing a Schur complement argument. This LMI, obtained systematically, offers a direct extension of the linear stability results [9, Thm. 3].
@@
where $Q:\mathbb{R}^{n}\rightarrow\mathbb{R}^{s-n}$ contains all the nonlinear functions in the library. The goal here is to stabilize the nonlinear system by rendering the closed-loop dynamics linear. In this case, the set of desired closed-loop systems is given by
$\mathcal{F}_{\rm des}:=\left\{\begin{bmatrix}\bar{F}&0\end{bmatrix}:\exists P>0,\quad\bar{F}^{T}P\bar{F}-P\leq 0\right\},$
@@
$\mathcal{F}_{\rm des}\cap\mathcal{F}_{\rm att}^{\rm data\,}=\left\{\begin{bmatrix}\bar{F}&0\end{bmatrix}\in\mathcal{F}_{\rm des}:\operatorname{im}\left[\begin{array}[]{@{}ll@{}}\bar{F}&0\\ \hline\cr\quad I_{s}\end{array}\right]\subseteq\operatorname{im}\begin{bmatrix}X_{1}\\ Z_{0}\end{bmatrix},\right\}.$
@@
A solution satisfying the above constraints exists under Assumption 1. By Theorem 3, given any feasible solution $(\bar{F},P,G)$ to the above constraints, the controller $u=KZ(x)$ with $K=U_{0}G$ solves Problem 1, namely the closed-loop system becomes a linear Schur stable system as desired. This coincides with the recent result [17, Thm. 1]. $\square$
Next, we move away from stabilization and discuss an example with a different control objective.
@@
$\mathcal{F}_{\rm des}=\left\{\begin{bmatrix}0&1&0&0^{T}_{s-3}\\ -\mu^{2}&1+\mu^{2}&-\frac{1}{3}\mu^{2}&0^{T}_{s-3}\end{bmatrix}:{\mu_{\ell}}\leq\mu\leq\mu_{u},\right\},$
@@
$X_{1}$ $\begin{bmatrix}g_{1}&g_{2}&g_{3}\end{bmatrix}=\begin{bmatrix}0&1&0\\ -\mu^{2}&1+{\mu^{2}}&-\frac{1}{3}{\mu^{2}}&\end{bmatrix},$
@@
III-B On the necessity of Assumption 1
As observed before, Problem 1 is solvable if and only if $\mathcal{F}_{\rm att}^{\rm model\,}\cap\mathcal{F}_{\rm des}\neq\emptyset$, where $\mathcal{F}_{\rm att}^{\rm model\,}$ is given by (3). The following result directly follows from Lemma 2.

### v2

# A Versatile Framework for Data-Driven Control of Nonlinear Systems

@@ Abstract
This note aims to provide a systematic investigation of direct data-driven control, enriching the existing literature not by adding another isolated result, but rather by offering a unifying, versatile, and broad framework that enables the generation of novel results in this domain. We formulate the nonlinear design problem from a high-level perspective as a set of desired controlled systems and propose systematic procedures to synthesize data-driven control algorithms that meet the specified design requirements. Various examples are presented to demonstrate the applicability of the proposed approach and its ability to derive new insights and results, illustrating the novel contributions enabled by the framework.
@@ I Introduction
The expanding interest in data-driven methodologies is fuelled by their potential to circumvent the limitations of model-based techniques, particularly in the realm of nonlinear systems where a priori given mathematical models may not be available. Nonlinear systems are characterized by their rich dynamics, which are not amenable to linear analysis and require a more involved approach to control design. As such, the field has seen diverse data-guided strategies being proposed and tested, ranging from virtual reference control [1], kernel-based design [2], intelligent PID [3] and sampled-data model-free control [4, 5] to nonlinear data-enabled model predictive control [6], [7], [8, Sec. 4], dynamic mode decomposition [9], Koopman design [10], first-order approximation [11, Subsection V.B] and Taylor-expansion [12, 13].
@@
The current work belongs to the category of direct data-driven control design aiming to reduce the task of controller design to programs that are stated in terms of data collected from the system. Due to its inherent complexity, the landscape of direct data-driven control for nonlinear systems is fragmented with solutions that often target specific problem settings or system types. Examples include stabilization of bilinear systems [14, 15], polynomial systems [16, 17], rational systems [18], and flat systems [19]. Such specialization is valuable as it provides concrete solutions to the control problems under investigation. Nonetheless, within the class of approaches that yield one-shot, data-dependent matrix inequalities for controller synthesis [20], there remains a need for a coherent framework that can address a broader spectrum of nonlinear control challenges. While the focus on particular instances or system models has driven substantial contributions, a unified approach could illuminate the underlying principles, fostering even greater advancements in the field.
@@
Recognizing the need and the benefits for such a unified approach, this note introduces a systematic framework for the direct data-driven design of nonlinear control systems. We begin by embedding the design specifications in a set of desired closed-loop systems. Subsequently, we establish systematic procedures to obtain the control gain from the system data such that the design specifications are enforced. The controller parametrization in [11], [21], serves as an inspiring element of the data-driven procedures laid out in this manuscript. The proposed approach distinguishes itself through its unifying, versatile, and innovative aspects, aiming not only to integrate several existing results into a cohesive framework but also to extend their applicability and enable the generation of new results in the domain of nonlinear direct data-driven control. Furthermore, the approach provides a systematic treatment of deriving control algorithms from data. This systematic procedure ensures that diverse objectives, such as stabilization, performance, and input-output properties can be seamlessly adapted and integrated into the design. Several cases are presented to highlight the versatility of the framework in systematically deriving new insights and results, showcasing its ability to enable novel contributions.
@@
The structure of the manuscript is as follows. The structure of the manuscript is as follows: The problem of interest is formulated in Section II, the solution is provided in Subsection III-A, and solvability conditions for the nonlinear control problem are discussed in Subsection III-B. Extensions of the results to open systems are provided in Subsection III-C. Finally, the manuscript closes with conclusions in Section IV.
@@ II Problem setup
$x^{+}=f(x)+Bu,$
where $x\in\mathbb{R}^{n}$, $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is locally Lipschitz, and $B\in\mathbb{R}^{n\times m}$. The notation $x^{+}$ denotes either the state variables shifted in time or the time derivative of the states depending on the setting (i.e. continuous or discrete time). We restrict our attention to nonlinear systems that are linear in the control input. While the map $f$ and the input matrix $B$ are assumed to be unknown, we assume that there exists a known library of nonlinear functions, denoted by $Z:\mathbb{R}^{n}\rightarrow\mathbb{R}^{s}$, such that $f$ can be written as $AZ(x)$ for some unknown matrix $A\in\mathbb{R}^{n\times s}$. Hence, the nonlinear system is represented by
@@
Under a nonlinear state feedback protocol $u=KZ(x)$, $K\in\mathbb{R}^{m\times s}$, the closed-loop system admits the form $x^{+}=F_{K}Z(x)$ with $F_{K}:=A+BK$, $F_{K}\in\mathbb{R}^{n\times s}$. The aim of the nonlinear controller is to impose some desired stability/performance properties on the closed-loop system. Rather than fixing a particular choice of a closed-loop, we consider a set of desired closed-loop dynamics. A set of desired closed-loop systems can be parametrized as
@@
where $\mathcal{V}$ denotes the set of Lyapunov functions satisfying $\alpha\left\|x\right\|^{2}\leq V(x)\leq\beta\left\|x\right\|^{2}$ for some $\alpha,\beta>0$. Restricting to quadratic Lyapunov functions, i.e. $V(x)=x^{T}Px$, the latter can be written more explicitly as
@@
Clearly, in the above examples, any $F^{\star}\in\mathcal{F}_{\rm des}$ is such that the closed-loop system (2) meets the design criterion. We also note that the Lyapunov dissipation inequalities in the presented examples of $\mathcal{F}_{\rm des}$ can be relaxed to a neighborhood around the equilibrium to capture local rather than global stabilization requirements.  $\square$
The central problem of this note is formulated as follows.
@@
$X_{1}=AZ_{0}+BU_{0}.$ (3)
The equality (3) is consistent with (1); specifically, the columns of $U_{0}$ corresponds to the input samples, those of $Z_{0}$ corresponds to the library $Z(\cdot)$ evaluated at the state samples, and the columns of $X_{1}$ are the state samples shifted in time in discrete time, or the time derivative of the state in continuous time. As will be observed, access to data samples compensates for the lack of knowledge in the system dynamics, enabling a data-driven solution to Problem 1.
@@ III-A Solutions to Problem 1
First, we emphasize that not any desired closed-loop system is attainable. Noting (1), a closed-loop system $x^{+}=FZ(x)$ is attainable if and only if there exists $K$ such that $A+BK=F$. Hence, the set of attainable, not necessarily desired, closed-loop systems can be parametrized by the following set:
We note that working with either the model or the data does not affect the choice or formulation of $\mathcal{F}_{\rm des}$. On the other hand, since we aim to use the collected data as a proxy for the model, we require a counterpart of (4) in terms of data. To this end, we define
@@
Proof. Item (i): If $\mathcal{F}_{\rm att}^{\rm data\,}$ is nonempty, then $\operatorname{im}I_{s}\subseteq\operatorname{im}Z_{0}$ and thus $Z_{0}$ has full row rank. Conversely, if $Z_{0}$ has full row rank, then $X_{1}Z_{0}^{+}\in\mathcal{F}_{\rm att}^{\rm data\,}$ for any right inverse $Z_{0}^{+}$ of $Z_{0}$.
Item (ii): Observe that the set $\mathcal{F}_{\rm att}^{\rm model\,}$ in (4) can be equivalently written as
@@
where we used (3) to write the first equality.
Item (iii): The inclusion in (7) can be replaced by equality if and only if the matrix $\left[\begin{smallmatrix}Z_{0}\\ U_{0}\end{smallmatrix}\right]$ has full row rank, where the “only if” part follows from the fact that the matrix $B$ and thus[$\left[\begin{smallmatrix}A&B\\ I_{s}&0\end{smallmatrix}\right]$ has full column rank. To conclude the proof, it remains to show that
$\mathcal{F}_{\rm att}^{\rm model\,}=\mathcal{F}_{\rm att}^{\rm data\,}\Longleftrightarrow\operatorname{im}\begin{bmatrix}X_{1}\\ Z_{0}\end{bmatrix}=\operatorname{im}{\begin{bmatrix}A&B\\ I_{s}&0\end{bmatrix}}$
The $\Leftarrow$’ direction trivially follows from (5) and (6). To prove the other direction, suppose $\mathcal{F}_{\rm att}^{\rm model\,}=\mathcal{F}_{\rm att}^{\rm data\,}$. Noting (7), it suffices to show that $\operatorname{im}[\begin{smallmatrix}A&B\\ I_{s}&0\end{smallmatrix}]\subseteq\operatorname{im}[\begin{smallmatrix}X_{1}\\ Z_{0}\end{smallmatrix}]$. Observe that for any $\Theta\in\mathbb{R}^{m\times s}$, the matrix $F:=A+B\Theta$ belongs to $\mathcal{F}_{\rm att}^{\rm model\,}$ and thus to $\mathcal{F}_{\rm att}^{\rm data\,}$. Hence, for any $\Theta\in\mathbb{R}^{m\times s}$, there exists a solution $\Xi$ to the following equation:
$\begin{bmatrix}A+B\Theta\\ I_{s}\end{bmatrix}=\begin{bmatrix}X_{1}\\ Z_{0}\end{bmatrix}\Xi.$
Now, we select two choices $\Theta=0$ and $\Theta=\bar{\Theta}$, for some $\bar{\Theta}\in\mathbb{R}^{m\times s}$ having full row rank. Then, there exist $\Xi$ and $\bar{\Xi}$ satisfying
$\begin{bmatrix}A&B\\ I_{s}&0\end{bmatrix}\begin{bmatrix}I_{s}&I_{s}\\ 0&\bar{\Theta}\end{bmatrix}=\begin{bmatrix}X_{1}\\ Z_{0}\end{bmatrix}\begin{bmatrix}\Xi&\bar{\Xi}\end{bmatrix}$.
By using the fact that $\begin{smallmatrix}I_{s}&I_{s}\\ 0&\bar{\Theta}\end{smallmatrix}]$ has full row rank, we conclude that $\operatorname{im}[\begin{smallmatrix}A&B\\ I_{s}&0\end{smallmatrix}]\subseteq\operatorname{im}[\begin{smallmatrix}X_{1}\\ Z_{0}\end{smallmatrix}]$, which completes the proof.
$\blacksquare$
Observe that Problem 1 is solvable if and only if $\mathcal{F}_{\rm att}^{\rm model\,}\cap\mathcal{F}_{\rm des}\neq\emptyset$; namely, the desired set $\mathcal{F}_{\rm des}$ and the attainable set $\mathcal{F}_{\rm att}^{\rm model\,}$ should at least share a common element. Motivated by the fact that we work directly with the data rather than the model, we replace $\mathcal{F}_{\rm att}^{\rm model\,}$ by its data-based subset $\mathcal{F}_{\rm att}^{\rm data\,}$ in the latter solvability condition, which results in the following assumption:
@@
Proof. The fact that the set $\mathcal{K}$ is nonempty follows from Assumption 1 and (5). Now, suppose $K\in\mathcal{K}$. Then, we find that
@@
where the second and last equality follow from (9), and the third one from (3). As $F^{\star}\in\mathcal{F}_{\rm des}$, we conclude that $(A+BK)\in\mathcal{F}_{\rm des}$ and the controller $KZ(x)$ solves Problem 1. $\blacksquare$
Theorem 3 provides a systematic procedure for solving Problem 1 using data:
@@
Next, we apply the above procedure to a series of illustrative examples that demonstrate its applicability and versatility across different scenarios.
@@
As the first example, we consider the control objective of rendering the equilibrium of a nonlinear system asymptotically stable by stabilizing the linearized dynamics. We note that the data is obtained from the original nonlinear system. It is easy to see that the set of desired closed-systems takes the form:
@@
where $Z^{\prime}$ is shorthand notation for the Jacobian of $Z$, and, for simplicity, the equilibrium is assumed to be at the origin. Following the second step of the procedure, we look for $F^{\star}\in\mathcal{F}_{\rm des}\cap\mathcal{F}_{\rm att}^{\rm data\,}$. This gives rise to the following set of constraints:
$FZ^{\prime}(0)P$ $(FZ^{\prime}(0))^{T}-P<0,\quad P>0,$ (11a)
$F$ $=X_{1}G,$ (11b)
$I_{s}$ $=Z_{0}G.$ (11c)
The aforementioned constraints can be transformed into a LMI through a suitable change of variables. Specifically, by defining $Y:=GZ^{\prime}(0)P$ and employing a Schur complement argument, we obtain
$\begin{bmatrix}P&X_{1}Y\\ Y^{T}X_{1}^{T}&P\end{bmatrix}>0,\quad Z^{\prime}(0)P-Z_{0}Y=0.$ (12)
If the LMI (12) admits a solution $(P,Y)$, then the constraints in (11) are satisfied for $(P,G)$ with
$G=Z_{0}^{+}+(I-Z_{0}^{+}Z_{0})YP^{-1}{Z^{\prime}(0)}^{\ell},$
as can be verified by direct substitution, where $Z_{0}^{+}:=Z_{0}^{T}(Z_{0}Z_{0}^{T})^{-1}$ and $Z^{\prime}(0)$$^{\ell}$ is any left inverse of $Z^{\prime}(0)$. A simple choice of this left inverse is obtained by partitioning the library as in (13) and setting ${Z^{\prime}(0)}^{\ell}=\begin{bmatrix}I_{n}&0\end{bmatrix}$. The resulting controller from the third step of the procedure is then given by $K=U_{0}G$. The LMI (12), obtained systematically, provides a generalization of the linear stabilization results [11, Thm. 3] to nonlinear systems.  $\square$
@@
where $Q:\mathbb{R}^{n}\rightarrow\mathbb{R}^{s-n}$ contains all the nonlinear functions in the library. The goal here is to stabilize the nonlinear system by rendering the closed-loop dynamics linear Schur stable. In this case, the set of desired closed-loop systems is given by
$\mathcal{F}_{\rm des}:=\left\{\begin{bmatrix}\bar{F}&0\end{bmatrix}:\exists P>0,\quad\bar{F}^{T}P\bar{F}-P<0\right\},$
@@
[…]
@@
A solution satisfying the above constraints exists under Assumption 1. By Theorem 3, given any feasible solution $(\bar{F},P,G)$ to the above constraints, the controller $u=KZ(x)$ with $K=U_{0}G$ solves Problem 1, namely the closed-loop system becomes a linear Schur stable system as desired. The result systematically obtained here coincides with the recent finding in [21, Thm. 1].
$\square$
Example 6 (Diagonal stabilization)
Consider a system in continuous-time with $Z(x)$ given by
$Z(x)=\left[\begin{smallmatrix}\phi_{1}(x_{1})\\ \phi_{2}(x_{2})\\ \ldots\\ \phi_{n}(x_{n})\\[2.84526pt] Q(x)\end{smallmatrix}\right],$
where $\phi_{i}:\mathbb{R}\rightarrow\mathbb{R}$ is a strictly monotone map for each $i$, $\phi_{i}(0)=0$, and $Q:\mathbb{R}^{n}\rightarrow\mathbb{R}^{s-n}$ contains additional nonlinear functions in the library. The desired closed-loop systems are given by
$\dot{x}=M\phi(x),$ (14)
where $\phi(x)=\operatorname{col}(\phi_{i}(x_{i}))$ and $M\in\mathbb{R}^{n\times n}$ is a diagonally stable matrix, i.e., $M^{T}D+DM<0$ for some diagonal positive definite matrix $D$. The motivation behind diagonal stability is as follows. Consider the Lyapunov function candidate $V(x):=\sum_{i=1}^{n}d_{i}\int_{0}^{x_{i}}\phi_{i}(y_{i})dy_{i}$, with $d_{i}$ being the $i$th diagonal element of $D$. Then, the time derivative of $V$ along the solutions of (14) is given by $\dot{V}(x)=\phi(x)^{T}DM\phi(x)\leq 0,$ where the inequality follows from diagonal stability property of $M$. Subsequently, asymptotic stability of the equilibrium follows from the implication $\phi(x)=0\Rightarrow x=0$. Now, consistent with (14), define
$\mathcal{F}_{\rm des}:=\left\{\begin{bmatrix}M&0\end{bmatrix}:\exists D=\operatorname{diag}(d_{i})>0,\,M^{T}D+DM<0\right\}.$
Following the proposed procedure, we search for an element in $\mathcal{F}_{\rm des}\cap\mathcal{F}_{\rm att}^{\rm data\,}$, which corresponds to finding $G_{1},G_{2},D$ satisfying
$M^{T}D+DM$ $<0,\quad D=\operatorname{diag}(d_{i})>0,$
$\bar{M}=$ $X_{1}G_{1},\quad 0=X_{1}G_{2},\quad I_{s}=Z_{0}\begin{bmatrix}G_{1}&G_{2}\end{bmatrix}$.
The corresponding controller is given by $u=U_{0}(G_{1}\phi(x)+G_{2}Q(x))$.  $\square$
We note that in data-driven stabilization procedures, what is utilized is not the specific Lyapunov function itself but rather a parametrized class of candidate functions. The design procedure searches within this class to identify a suitable Lyapunov function that certifies the stability of the closed-loop system. Next, we move away from stabilization and discuss an example with a different control objective.
@@
$\mathcal{F}_{\rm des}=\left\{\begin{bmatrix}{1}&1&0&0^{T}_{s-3}\\ -\mu^{2}&1+\mu^{2}&-\frac{1}{3}\mu^{2}&0^{T}_{s-3}\end{bmatrix}:{\mu_{\ell}}\leq\mu\leq\mu_{u},\right\},$
@@
$X_{1}$ $\begin{bmatrix}g_{1}&g_{2}&g_{3}\end{bmatrix}=\begin{bmatrix}{1}&1&0\\ -\mu^{2}&1+{\mu^{2}}&-\frac{1}{3}{\mu^{2}}&\end{bmatrix}$,
@@
III-B On the solvability of Problem 1 from data
In this subsection, we demonstrate that the condition $\mathcal{F}_{\rm att}^{\rm data\,}\cap\mathcal{F}_{\rm des}\neq\emptyset$, as assumed in Assumption 1 and with $\mathcal{F}_{\rm att}^{\rm data\,}$ defined in (5), is generally non-conservative when solving Problem 1 from data.
First, recall that Problem 1 is solvable if and only if $\mathcal{F}_{\rm att}^{\rm model\,}\cap\mathcal{F}_{\rm des}\neq\emptyset$, where $\mathcal{F}_{\rm att}^{\rm model\,}$ is given by (4). The following result directly follows from Lemma 2.